Sin raise to 4 theta - 2sin square theta = 00 Find all the angles lying between 0 degree and 360 degree

Grade:11

1 Answers

Jitender Singh IIT Delhi
9 years ago
Let theta = x
sin^4 x – 2sin^2 x = 0
sin^2 x(sin^2 x – 2) = 0
We have 2 equations
sin^2 x = 0.......(1) sin^2 x = 2........(2)
From (1), we have From (2), we have
sin x = 0 sin x = ^{}^{^{}}sqroot(2) sin x = -sqroot(2)
x = nπ Not possible (because -1 <= sin x <= 1)
So final solution is x = 0,π, 2π
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty